OpenStudy (anonymous):

log8(x^2-30)=log8(x)

OpenStudy (anonymous):

as there are logs both sides: x^2 - 30 = x solve this quadratic and you're home

OpenStudy (campbell_st):

equate the terms since they are logs of the same base x^2 - 30 = x x^2 - x - 30 = 0 (x - 6)(x+5) = 0 solve for x

OpenStudy (anonymous):

Make sure x value is positive!

OpenStudy (anonymous):

\[\LARGE \log_8(x^2-30)=\log_8(x) \quad \iff x^2-30=x\] \[\LARGE x^2-x-30=0 \] use quadratic formula , \[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] or you can factor it out, if you want it's up to you :)

OpenStudy (sburchette):

You can subtract log8(x) from both sides and get\[\log_8(x^2-30)-\log_8(x)=0\]Using the subtraction property for logs we get\[\log_8((x^2-30)/x)=0\]Going to exponential form you get \[8^0=(x^2-30)/x\]\[x=x^2-30\]\[x^2-x-30=0\]\[(x-6)(x+5)=0\]\[x=6, x=-5\] Because a logarithm is undefined for negative values of x, only x=6 is valid.

OpenStudy (anonymous):

Thank you everyone!